OpenStudy (anonymous):

Need help solving: log2(y+2)-1=log2(y-2)

OpenStudy (anonymous):

the 2 being the base

OpenStudy (anonymous):

2(y-2) = (y+2) 2y -y = 2 + 4 -> y = 6

OpenStudy (anonymous):

chlorophyll i dont follow

OpenStudy (anonymous):

1 = log2(2) => left hand side is log2(y+2) - log2(2) log laws state that this can be rewritten as log2((y+2)/2) so the equation reads log2((y+2)/2)=log(y-2) raising both 'onto' a power of 2 yields (y+2)/2 = y-2 multiply by 2 y+2 = 2y - 4 rearrange for y; 6 = y

OpenStudy (anonymous):

Awesome thank you!